What is the equation of the asymptote of the hyperbola x^2 - y^2 = 1?
Answer by ikleyn(52802)

.

To get the asymptotes equations, REPLACE the "1" in the right side by "0"


    x^2 - y^2 = 0


and factor the left side


    (x-y)*(x+y) = 0.


The equations  

    x - y = 0

and

    x + y = 0

are the equation's of the asymptotes, in this case.

Solved, answered and explained.